OpenStudy (anonymous):

any one know anything about the terms i? i have a problem i cant figure out. -(sqrt)-496(both are negatives)

OpenStudy (anonymous):

So you have: \[-\sqrt{-496}\] Well, i is defined as \[\sqrt{-1}\]. So you can rewrite this as: \[-\sqrt{(-1)(496)}\]. Then: \[-(\sqrt{-1}\sqrt{496})\]. So this becomes. \[-\sqrt{496}i\]

OpenStudy (anonymous):

Does this clear it up?

OpenStudy (anonymous):

is that the answer? im just not sure really how this i thing works...kinda confusing...or atleast math in general confuses me..

OpenStudy (anonymous):

is that in its lowest terms?

OpenStudy (anonymous):

That is "exact" answer. Yes. sqrt(496) is a decimal. (not a nice one xP)

OpenStudy (anonymous):

If you need more help/explanation just tell me :)

OpenStudy (anonymous):

lol thanks